# Defects and Impurity Properties of VN precipitates in ARAFM Steels: Modelling using a Universal Machine Learning Potential and Experimental Validation

### v1

@@ 4 Discussion
APT results show that VN precipitates may not be modelled appropriately as only a V-N system and that ternary or more complex systems may be required. However, DFT, whilst commonly used for its accuracy, is too computationally expensive to study such systems. Here, the Orb UMLIP was employed to explore a larger series of systems that cannot easily be explored with DFT alone, and it can do so capably. The UMLIP was able to approach DFT-level results for stable VN structures [10], including finding previously reported stable vacancy layers.
[…]
Both this work and prior work in the literature[28] indicate that VN accommodate N vacancies, with highly ordered layered vacancy structures as ground states. However, irradiation damage at 10s of dpa is expected to destroy long-range ordering in these precipitates. The particular temperature and fluence experienced by the material could result in a dynamic equilibrium by which irradiation destroys ordered structures generating randomly ordered vacancies, and a thermodynamic driving force drives diffusion to reorder the structure into vacancy layers.
In addition, the ternary systems studied suggest possible mechanisms of dissolution of VN. Due to irradiation damage, Fe is likely to be present in the VN precipitates, either via ballistic implantation, or by vacancy-driven migration. However, our calculations (Fig. 3) indicate that Fe may act to destabilise the VN.
@@ 5 Conclusions
A combined experimental and computational approach have led to both a better understanding of the VN precipitates in ARAFM steel and the effect of common irradiation defect on precipitate stability. These findings can be summarized as follows.
@@
    UMLIP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured over disordered point vacancies.
@@
    UMLIP calculations predict Fe substitutions to be unstable in VN.
@@
    Switching from bulk to local reference states changed the apparent stability of vacancies in VN. This highlights how a calculated binary phase diagram may lead to inaccurate expectations when predicting phase stability of precipitates in more complex alloy systems, such as MX precipitates in ARAFM steel. Reconsidering the reference states is a simple way to gauge phase stability in these alloys. Further calculations of the phase stabliity of other ARAFM steel MX phase precipitates with respect to local reference states are needed.

### v2

@@ 4 Discussion
APT results show that VN precipitates in these ARAFM may not be modelled appropriately as only a V-N system and that ternary or more complex systems may be required. However, DFT, whilst commonly used for its accuracy, is too computationally expensive to study such systems. Here, the NEP89 potential was finetuned and employed to explore a larger systems that cannot easily be explored with DFT alone, and it can do so capably. The potential was able to replicate DFT-level results for stable VN structures [32], including finding previously reported stable vacancy layers.
Experimentally, the pre-irradiation vacancy and solute concentration for the VN precipitates is difficult to measure due to the small precipitate size and the limitations of APT. The point defect calculations for VN show a preference for vacancy-driven non-stoichiometry with a thermodynamic preference for layered vacancy structures proposed by [14, 13]. However, when accounting for the local environment by adjusting the reference states the preference for N vacancies was weaker than previously reported reducing the range of stable stoichiometries present, highlighting a need to reconsider MX-type precipitates in the context of the surrounding steel matrix. The low vacancy concentration structures (5% to 50%) were still present on the convex hull. Including N vacancies and Cr-V substitutions may partially explain the small experimental lattice parameter of 4.035 Å, although the Cr had little effect.
Both this work and prior work in the literature[31] indicate that VN can accommodate N vacancies, with highly ordered layered vacancy structures as ground states. However, irradiation damage at 10s of dpa is expected to destroy long-range ordering in these precipitates. The particular temperature and fluence experienced by the material could result in a dynamic equilibrium by which irradiation destroys ordered structures generating randomly ordered vacancies, and a thermodynamic driving force drives diffusion to reorder the structure into vacancy layers. It is shown here that there remains ordering of vacancies at the target operating temperature of over 800 K, albeit the ordering is reduced. In addition, the effect of single cascades suggests that low to moderate vacancy concentrations of (5% to 50%) may be more robust to irradiation damage on account of deposited incoming energy leading to the reordering of defects in addition to creating them. Lastly, the effect of solutes seems to play a major role in irradiation damage of VN, and as such caution should considered when such solutes are excluded. In particular, solutes act to disrupt the vacancy ordering on account of introducing strain into the lattice. Such strain leads to the formation of solute plus vacancy clusters and is shown with cascades (Fig. 6) to increase the stored energy relative to the 5% vacancy condition. Whilst the lattice Monte Carlo model showed vacancy-solute cluster formation, it is important to recognise the model does not account for kinetics. It is known that cluster composition and formation depends partially on the diffusion rates of solutes and defects. In particular, it is found here that W is a key component of solute clustering in VN, although given its large atomic radii it is expected to diffuse slowly. A model accounting for both the energetics and kinetics may not find the same vacancy behaviour.
@@ 5 Conclusions
A combined experimental and computational approach have led to both a better understanding of the VN precipitates in ARAFM steel and the effect of common irradiation defects on precipitate stability. These findings can be summarized as follows.
@@
    NEP and DFT results both predict sub-stoichiometric ground states in the VN system with ordered vacancy layers favoured over disordered point vacancies.
@@
    Switching from bulk to local reference states changed the apparent stability of vacancies in VN. This highlights how a calculated binary phase diagram may lead to inaccurate expectations when predicting phase stability of precipitates in more complex alloy systems, such as MX precipitates in ARAFM steel. Reconsidering the reference states is a simple way to gauge phase stability in these alloys. Further calculations of the phase stability of other ARAFM steel MX phase precipitates with respect to local reference states are needed.
@@
    Lattice based Monte Carlo confirmed the presence of vacancy ordering at operational temperatures of 800-900+ K in ARAFM steel.
  • 6.
    PKA collision cascades showed the importance of vacancies and solutes in stored defect energy introduced from cascades. In particular, the 5, 14 and 26% vacancy concentrations mitigated damage, whilst the addition of solutes reversed this for the 5% solute case tested.
6 CRediT authorship contribution statement
R. S. Stroud: Conceptualization, data curation, investigation, formal analysis, writing—original draft, and visualization. C. Reynolds: Writing—review and editing and supervision. T. Melichar: Data curation, investigation, formal analysis and visualization. J. Haley: Investigation and writing—review and editing. M. Carter: investigation. M. Moody: Funding acquisition. C. Hardie: Supervision, writing—review and editing. D. Bowden: Supervision. D. Nguyen-Manh: Conceptualization, formal analysis, writing—review and editing and supervision. M. R. Wenman: Conceptualization, formal analysis, writing—review and editing, supervision, and funding acquisition
@@ Acknowledgements
Supplementary Information
Figure 7: Density functional theory convergence testing for kpoint spacing at 520 eV energy cutoff.
Figure 8: Convergence testing for energy cutoff at a kpoint spacing of 0.2 Å-1.
Figure 9: The testing dataset parity plots for the finetuned NEP89 potential against the DFT values for a) energy, b) force, c) stress, and d) virial.